Exercise 15-14 Factory overhead computed, applied, and adjusted LO P3, P4 The following information applies to the questions displayed below In December 2014, Custom Mfg established its predetermined overhead rate for jobs produced during year 2015 by using the following cost predictions: overhead costs, $1,080,000, and direct labor costs, $600,000 At year-end 2015, the companys records show that actual overhead costs for the year are $891,400. Actual direct labor cost had been assigned to jobs as follows. Jobs completed and sold Jobs In finished goods inventory Jobs in work in process inventory 360,000 81,000 49,000 Total actual direct labor cost $490,000
media%2F7fc%2F7fcafdaa-d944-4a40-93e5-ff
media%2F6fc%2F6fc4e97f-0314-4485-902f-84
media%2Fe76%2Fe76bfd8a-439f-4379-a7e0-be
Answer #1

1

Estimated overhead costs/Estimated direct labor costs = Overhead rate

1080000/600000 = 180%

2&3

Factory overhead

Actual overhead

$891400

$882000

Applied overhead

Under applied overhead(891400-882000)

$9400

Overhead applied to jobs = 490000*180% = $882000

4

Cost of goods sold

$9400

        Factory overhead

$9400